Problem 5-18 Journal Entries; T-Accounts; Cost Flows [LO4, LO5, LO7]

Ravsten Company uses a job-order costing system. On January 1, the beginning of the current year, the company’s inventory balances were as follows:

  
  Raw materials $ 16,500
  Work in process $ 10,200
  Finished goods $ 30,100

The company applies overhead cost to jobs on the basis of machine-hours. For the current year, the company estimated that it would work 36,100 machine-hours and incur $155,230 in manufacturing overhead cost. The following transactions were recorded for the year:

  1. Raw materials were purchased on account: $202,000.
  2. Raw materials were requisitioned for use in production: $191,000 (80% direct and 20% indirect).
  3. The following costs were incurred for employee services:
  
  Direct labour $ 160,800
  Indirect labour $ 27,200
  Sales commissions $ 36,300
  Administrative salaries $ 80,400
  1. Heat, power, and water costs were incurred in the factory: $42,450.
  2. Prepaid insurance expired during the year: $10,500 (85% relates to factory operations, and 15% relates to selling and administrative activities).
  3. Advertising costs were incurred, $50,500.
  4. Depreciation was recorded for the year: $60,600 (90% relates to factory operations, and 10% relates to selling and administrative activities).
  5. Manufacturing overhead cost was applied to production. The company recorded 40,200 machine-hours for the year.
  6. Goods that cost $484,400 to manufacture according to their job cost sheets were transferred to the finished goods warehouse.
  7. Sales for the year totalled $704,700 and were all on account. The total cost to manufacture these goods according to their job cost sheets was $479,200.

Required:

3-a. Is manufacturing overhead underapplied or overapplied for the year?

  • Underapplied overhead

  • Overapplied overhead

3-b. Prepare a journal entry to properly dispose of any balance in the Manufacturing Overhead account. (Do not round intermediate calculations and round your final answers to 2 decimal places. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

4. Prepare an income statement for the year. (Round intermediate calculations to nearest whole number.)
Answer #1

3a) Applied overhead = 155230/36100*40200 = 172860

Actual overhead = (191000*20%+27200+42450+10500*85%+60600*90%) = 171315

Over applied overhead

3b) Journal entry

Date account and explanation Debit Credit
Manufacturing overhead 1545
Cost of goods sold 1545
(To record over applied overhead)

4) Income statement

Sales 704700
Cost of goods sold (479200-1545) 477655
Gross profit 227045
Operating expense
Sales commissions 36300
Administrative salaries 80400
Advertising expense 50500
Insurance expense 1575
Depreciation 6060 174835
Net income 52210